# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ Appendix D Stationary deme occupancy
This mean-field derivation of $\Omega_{\rm occ}$ is accurate for large clique metapopulations but, as it ignores spatial correlations, it is a crude approximation for cycles and grids, see Fig. 9. In particular, $\Omega_{\rm occ}$ overestimates the number of occupied demes in cycles when $\psi$ is not much larger than 1. However, $\psi=mK\tau_{E}$ allows us to distinguish between different regimes and provides a sound estimate of the number of occupied demes in the intermediate regime when $\psi(m,K)\approx me^{K}$ is sufficiently bigger than $1$.
@@ Appendix E Average number of active edges on the square grid and the influence of the spatial structure
In principle, the coarse-grained description of the $M/W$ competition holds for any regular circulation graph. However, this approach requires the determining of the number of active edges, which, except for cycles (one dimension) and cliques, is a difficult task due to complex spatial correlations between demes. Here, we consider the case of the square grid (with periodic boundaries), and illustrate how to estimate the average number of active edges when the metapopulation consists of only one single fully-occupied $M$ deme and all other demes are occupied by $W$ cells.
@@
Figure 10: Near-optimal condition for the idealised treatment on a cycle and grid metapopulation. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for the cycle (a-d) and grid (e-h) metapopulations, see Appendix E. Whitespace in panels (c) and (g) indicate the region of the parameter where at least one realisation for those parameters did not reach extinction by $t=10^{5}$. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (13). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Other parameters are $\Omega=16$, $m=10^{-4}$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
In the case of a large metapopulation, $\Omega\gg 1$, with unit spacing between neighbours, we assume that the mutant spreads outwards from the initial $M$ deme approximately forming a $M$-cluster having a circular front. If this circular $M$-cluster has a radius $r$, it has an area $\pi r^{2}$ containing a number $i$ of $M$ demes. The boundary of the circular $M$-cluster is of length $2\pi r$. Assuming that this length is equal to the number of $M$ demes on the boundary, we find that $r=\sqrt{i/\pi}$ and there are $2\sqrt{i\pi}$ boundary demes given $i$ demes of type $M$. We therefore estimate that the average number active edges for a grid is $E_{\rm grid}(i)\approx 2\sqrt{\pi i}$. We have notably used this approximation in the transitions rates Eq. (10) and Eq. (19) for the coarse-grained description of $M/W$ competition in static and time-varying environments in the regime of weak bottlenecks.
@@
In Figs. 2(a,c) and 4(c), we have found that the spatial structure has a barely noticeable influence on the fixation probability and mean fixation time when the carrying capacity is constant and in time-varying environments in the regime of weak bottlenecks. Fig. 10 shows the heatmaps on a cycle and a grid metapopulation for the “idealised treatment” proposed in Sec. IV.0.2, which are almost identical. This is in accord with Eqs. (25) predicting that the same migration rate yields the same near optimal conditions for the heatmaps of metapopulation on any regular graph, here a cycle and a grid. Simulation results confirm spatial structure is only responsible for minor quantitative changes in the region of the heatmaps corresponding to the near-optimal “treatment conditions”. This stems from the removal scenario characterising the idealised treatment being due to deme extinction which is mostly independent of ${\rm G}$ and $m$.
@@ Appendix F Further details on the intermediate dynamics in static environments
Figure 11: $\phi_{\rm int}$ vs. $K$ for $s=0.1$ (blue) and $s=0.01$ (red) on clique (solid lines / crosses), cycle (dashed lines / circles), and a grid (dotted lines / triangles). The different symbols and lines are almost indistinguishable. $\theta_{\rm int}$ vs. $K$ for the same parameters for a clique metapopulation. The vertical solid and dashed line indicate where the number of occupied demes $\Omega_{\text{occ}}\in[1,\Omega-1]$, see text. The dotted line indicates where $\psi=\Omega K$, i.e. every individual migrates in the time required for an independent deme extinction. Markers show simulation results, solid lines are predictions of Eq. (44), and dash-dot lines are predictions of Eq. (12). Other parameters are $m=10^{-4}$ and $\Omega=16$.
@@
In this section, we provide further details about the analysis of the intermediate dynamics in static environments, see Sec. III.0.3. Here, we complete the result given in the main text by considering metapopulation intermediate dynamics on a regular graph ${\rm G}$ (${\rm G}$-graph), obtaining the explicit results $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ for ${\rm G}=\{{\rm clique,cycle,grid}\}$ reported in Fig. 11 when the metapopulation initially consists of a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
The intermediate regime is characterised by $M/W$ competition and deme extinction. Therefore, in addition to invasions, a $W$ deme may become an $M$ deme through extinction followed by a recolonisation, i.e. $W\rightarrow\emptyset\rightarrow M$, where $\emptyset$ indicates an extinct deme. Similarly, an $M$-deme can be changed into a $W$-deme via $M\rightarrow\emptyset\rightarrow W$. We assume that there is initially a single $M$ deme in the metapopulation (and $\Omega-1$ demes of type $W$). With a probability $p_{\text{surv}}$ (see below), the initial $M$ deme survives to reach the dynamical equilibrium that consists of an unbreakable $M$-cluster whose size (number of $M$ demes) $i=0,1,\dots,\Omega_{\rm occ}$ grows and shrinks through invasions and extinction-recolonisation events. We assume that immediately $\Omega_{E}=\Omega/\psi$ demes go extinct, so that the metapopulation quickly reaches its equilibrium occupancy $\Omega_{\text{occ}}=\Omega\left(1-1/\psi\right)$. In this dynamical equilibrium, a $W$ deme can become an $M$ deme via $W\rightarrow\emptyset$ ($W$ deme extinction) at rate $r_{{\rm ext},W}$ followed by $\emptyset\rightarrow M$ (recolonisation by $M$) at rate $r_{{\rm rec},M}^{\rm G}$. The overall extinction-recolonisation reaction $W\rightarrow\emptyset\rightarrow M$ thus occurs at rate $1/(1/r_{{\rm ext},W}+1/r_{{\rm rec},M}^{\rm G})$. Here, the rate of $W$ deme extinction is $r_{{\rm ext},W}=(\Omega_{\text{occ}}-i)/\tau_{E}$ and $\tau_{E}$ is given by Eq. (6). We proceed similarly for the extinction of an $M$ deme and its recolonisation into a $W$ site according to $M\rightarrow\emptyset\rightarrow W$. Taking also into account the rate of invasion, see (10), the size $i$ of the $M$-cluster on a ${\rm G}$-regular graph varies according to the transition rates
$\widetilde{T}_{i}^{+}(m,{\rm G},K)$ $=mK\frac{E_{\rm G}(i)}{q_{\rm G}}\left[\rho_{M}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{E_{\rm G}(i)}\right],$ (42)
$\widetilde{T}_{i}^{-}(m,{\rm G},K)$ $=mK\frac{E_{\rm G}(i)}{q_{\rm G}}\left[\rho_{W}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{E_{\rm G}(i)}\right].$
With these rates, we can solve the following first-step analysis equations for the probability $\phi_{{\rm int},i}^{\rm G}$ that the dynamical equilibrium comprising an initial $M$-cluster of size $i$ contains only of occupied $M$ demes after a mean time $\theta_{{\rm int},i}^{\rm G}$:
@@
These equations are subject to the boundary conditions $\phi_{{\rm int},0}^{\rm G}=0,\phi_{{\rm int},\Omega_{\rm occ}}^{\rm G}=1$ and $\theta_{{\rm int},0}^{\rm G}=\theta_{{\rm int},\Omega_{\rm occ}}^{\rm G}=0$. We thus have $\phi_{\rm int}^{\rm G}\equiv p_{\text{surv}}\phi_{{\rm int},1}^{\rm G}$ and $\theta_{\rm int}^{\rm G}\equiv p_{\text{surv}}\theta_{{\rm int},1}^{\rm G}+(1-p_{\text{surv}})\tau_{E}$. The factor $p_{\text{surv}}=\frac{\Gamma(\Omega)}{\Gamma(\Omega_{\text{occ}})}\frac{1}{\Omega^{\Omega_{\text{occ}}}}$ is the probability that the initial $M$ deme reaches the dynamical equilibrium, while the contribution $(1-p_{\text{surv}})\tau_{E}$ to $\theta_{\rm int}^{\rm G}$ accounts for the probability that the initial $M$ deme goes extinct in a mean time $\tau_{E}$ (given by Eq. (6)) before reaching the equilibrium. The final expressions of $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ thus read
@@
$\widetilde{\gamma}_{\rm G}(i)\equiv\frac{\rho_{W}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{E_{\rm G}(i)}}{\rho_{M}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{E_{\rm G}(i)}},$ (45)
@@
We notice that for all graphs ${\rm G}$, the expressions of Eq. (44) coincide with those of Eq. (12) of the competition-dominated regime, with $\widetilde{\gamma}_{G}(i)\stackrel{{\scriptstyle\psi\gg 1}}{{\longrightarrow}}\gamma=\rho_{W}/\rho_{M}$. In Fig. 11, we find that the predictions of Eq. (44) are in good agreement with simulation results for all spatial structures ${\rm G}$. Moreover, we notice that the spatial dependence of $\phi^{\rm G}_{\text{int}}$ and $\theta^{\rm G}_{\text{int}}$ is barely noticeably.
@@ Appendix G Fixation probability in time-switching environments under weak bottlenecks
In static environments, where $K$ is constant, a generalisation of the circulation theorem guarantees that the fixation probability is independent on the migration rate and spatial structure of the metapopulation arranged on a circulation graph, see Eq. (12). This results from a correspondence between the fixation probability and the number of $M$ demes performing a biased random walk on $[0,\Omega]$ with a bias that is independent of $m$ and ${\rm G}$ [41].
In time-switching environments under weak bottlenecks (deme extinction is negligible) the correspondence is between the fixation $\Phi^{\rm G}(\nu,m)$ and the random walk (with absorbing boundaries) on $[0,\Omega]\times\{-1,1\}$ for the number of fully mutant demes in the environmental state $\xi=\pm 1$. As a consequence, $\Phi^{\rm G}(\nu,m)$ is the probability of absorption in the state $\Omega$. In this setting, defining the state of the random walk by $(i,\xi)$, where $i=0,1,\dots,\Omega$, the random walk moves to the right ($i\to i+1$) with a probability $r(i,\xi)$, to the left ($i\to i-1$) with a probability $\ell(i,\xi)$, or switches environment ($\xi\to-\xi$) with probability $\epsilon(i,\xi)$, where
@@
$\Phi^{\rm G}(\nu,m)$ thus coincides with the probability that the random walk defined by Eq. (47) gets absorbed in the state $i=\Omega$. For the fixation probability to be independent of $m$ and ${\rm G}$ then requires each of $r(i,\xi)$, $\ell(i,\xi)$, and $\epsilon(i,\xi)$ to be independent of $m$ and $\rm G$. However, this is generally not the case in the presence of environmental switching ($\nu>0$) due to the explicit and implicit dependence of Eq. (47) on $\nu$ (in the denominator and via $\mathcal{N}$ and $\rho_{M/W}$). This implies that in time-fluctuating environments the fixation probability $\Phi^{\rm G}(\nu,m)$ is expected to depend on $m$ and ${\rm G}$.
Interestingly however, Fig. 4(c) illustrates the almost unnoticeable dependence of $\Phi^{\rm G}(\nu,m)$ on the specific spatial structure. This is due to the overall similar impact of the factor $E_{\rm G}(i)/q_{G}$ for the various graphs. While differences arise when $m$ varies at fixed $\nu$ due to large variations in the timescales of the competition dynamics (significantly more invasions when $K=K_{+}$ than when $K=K_{-}$), varying spatial structure produces small changes in these timescales, and as such leads to only unnoticeable changes in $\Phi$.

### v2

@@ Appendix D Stationary deme occupancy
This mean-field derivation of $\Omega_{\rm occ}$ is accurate for large clique metapopulations but, as it ignores spatial correlations, it is a crude approximation for cycles and grids; see Fig. 9. In particular, $\Omega_{\rm occ}$ overestimates the number of occupied demes in cycles when $\psi$ is not much larger than 1. However, $\psi=mK\tau_{E}$ allows us to distinguish between different regimes and provides a sound estimate of the number of occupied demes in the intermediate regime when $\psi(m,K)\approx me^{K}$ is sufficiently bigger than $1$.
@@ Appendix E Average number of active edges on the square grid and the influence of the spatial structure
In principle, the coarse-grained description of the $M/W$ competition holds for any regular circulation graph. However, this approach requires the number of active edges for a given number of mutant demes to be known, which, except for cycles (one dimension) and cliques, is a difficult task due to complex spatial correlations between demes. Here, we consider the case of the square grid (with periodic boundaries), and illustrate how to estimate the average number of active edges when the metapopulation consists of only one single fully-occupied $M$ deme and all other demes are occupied by $W$ cells.
@@
Figure 10: Near-optimal condition for the idealised treatment strategy on a cycle and grid metapopulation. $(\nu,K_{-})$ heatmaps of $\Phi$, $\Theta$, $\Theta_{E}$ and $(1-\Phi)(1-\Theta/\Theta_{E})$ for the cycle (a-d) and grid (e-h) metapopulations; see Appendix E. White space in panels (c) and (g) indicates the parameter region where at least one realisation for those parameters did not reach extinction by $t=10^{5}$. Grey lines in panels (d) and (h) show the near-optimal conditions for the idealised treatment strategy: $\psi(m,K_{-})<1$ below the top horizontal line, $mK_{+}\theta_{E}>1$ above the bottom horizontal line, and $\nu\theta_{E}>1$ above the curved line, while the vertical line indicates where $\nu<1$ and $\theta_{E}$ from Eq. (15). The near-optimal treatment conditions is the yellowish cloud at the centre of the area enclosed by these lines. Other parameters are $\Omega=16$, $m=10^{-4}$, $s=0.1$, and $K_{+}=200$. In all panels, initially there is a single $M$ deme and $\Omega-1$ demes occupied by $W$.
@@
In the case of a large metapopulation, $\Omega\gg 1$, with unit spacing between neighbours, we assume that the mutant spreads outwards from the initial $M$ deme approximately forming an $M$-cluster with a circular front. If this circular $M$-cluster has a radius $r$, it has an area $\pi r^{2}$ containing a number $i$ of $M$ demes. The boundary of the circular $M$-cluster is of length $2\pi r$. Assuming that this length is equal to the number of $M$ demes on the boundary, we find that $r=\sqrt{i/\pi}$ and there are $2\sqrt{i\pi}$ boundary demes given $i$ demes of type $M$. We therefore estimate that the average number active edges for a grid is $E_{\rm grid}(i)\approx 2\sqrt{\pi i}$. We have notably used this approximation in the transitions rates Eq. (11) and Eq. (21) for the coarse-grained description of $M/W$ competition in static and time-varying environments in the regime of weak bottlenecks.
@@
In Figs. 2(a,c) and 4(c), we have found that the spatial structure has a barely noticeable influence on the fixation probability and mean fixation time when the carrying capacity is constant and in time-varying environments in the regime of weak bottlenecks. Fig. 10 shows the heatmaps on a cycle and a grid metapopulation for the “idealised treatment strategy” proposed in Sec. IV.2, which are almost identical. This is in accord with Eqs. (27) predicting that the same migration rate yields the same near optimal conditions for the heatmaps of metapopulation on any regular graph, here a cycle and a grid. Simulation results confirm spatial structure is only responsible for minor quantitative changes in the region of the heatmaps corresponding to the near-optimal “treatment conditions”. This stems from the removal scenario characterising the idealised treatment strategy being due to deme extinction which is mostly independent of $\rm G$} and $m$.
@@ Appendix F Further details on the intermediate dynamics in static environments
Figure 11: $\phi_{\rm int}$ vs. $K$ for $s=0.1$ (blue) and $s=0.01$ (red) on clique (solid lines / crosses), cycle (dashed lines / circles), and a grid (dotted lines / triangles). The different symbols and lines are almost indistinguishable. $\theta_{\rm int}$ vs. $K$ for the same parameters for a clique metapopulation. The vertical solid and dashed line indicate where the number of occupied demes $\Omega_{\text{occ}}\in[1,\Omega-1]$; see text. The dotted line indicates where $\psi=\Omega K$, i.e. every individual migrates in the time required for an independent deme extinction. Markers show simulation results, solid lines are predictions of Eq. (44), and dash-dot lines are predictions of Eq. (14). Other parameters are $m=10^{-4}$ and $\Omega=16$.
@@
In this section, we provide further details about the analysis of the intermediate dynamics in static environments; see Sec. III.3. Here, we complete the result given in the main text by considering metapopulation intermediate dynamics on a regular graph $\rm G$} ($\rm G$}-graph), obtaining the explicit results $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ for ${\rm G}=\{{\rm clique,cycle,grid}\}$ reported in Fig. 11 when the metapopulation initially consists of a single $M$ deme and $\Omega-1$ demes occupied by $W$. This region is concretely defined for values of $\psi$ such that $1<\psi<K\Omega$ for $\Omega\gg 1$, where the lower bound ensures $\Omega_{\text{occ}}=1$ from Eq. (9) and the upper bound is obtained by considering the limiting case where every individual is expected to migrate on the timescale of deme extinction. Using that $\psi\approx me^{K}$ for $K\gg 1$, we obtain the condition $\ln(1/m)\lesssim K\lesssim\ln(\Omega K/m)$.
@@
The intermediate regime is characterised by $M/W$ competition and deme extinction. Therefore, in addition to invasions, a $W$ deme may become an $M$ deme through extinction followed by a recolonisation, i.e. $W\rightarrow\emptyset\rightarrow M$, where $\emptyset$ indicates an extinct deme. Similarly, an $M$-deme can be changed into a $W$-deme via $M\rightarrow\emptyset\rightarrow W$. We assume that there is initially a single $M$ deme in the metapopulation (and $\Omega-1$ demes of type $W$). With a probability $p_{\text{surv}}$ (see below), the initial $M$ deme survives the short transient as the metapopulation reaches the dynamical equilibrium that consists of the $M$ deme, the remaining $W$ demes, and extinct demes. The number of $M$ demes $i=0,1,\dots,\Omega_{\rm occ}$ grows and shrinks through invasions and extinction-recolonisation events. We assume that immediately $\Omega_{E}=\Omega/\psi$ demes go extinct, so that the metapopulation quickly reaches its equilibrium occupancy $\Omega_{\text{occ}}=\Omega\left(1-1/\psi\right)$. In this dynamical equilibrium, a $W$ deme can become an $M$ deme via $W\rightarrow\emptyset$ ($W$ deme extinction) at rate $r_{{\rm ext},W}$ followed by $\emptyset\rightarrow M$ (recolonisation by $M$) at rate $r_{{\rm rec},M}^{\rm G}$. The overall extinction-recolonisation reaction $W\rightarrow\emptyset\rightarrow M$ thus occurs at rate $1/(1/r_{{\rm ext},W}+1/r_{{\rm rec},M}^{\rm G})$. Here, the rate of $W$ deme extinction is $r_{{\rm ext},W}=(\Omega_{\text{occ}}-i)/\tau_{E}$ and $\tau_{E}$ is given by Eq. (7). We proceed similarly for the extinction of an $M$ deme and its recolonisation into a $W$ site according to $M\rightarrow\emptyset\rightarrow W$. Taking also into account the rate of invasion (see Eq. (11)) the number of $M$ demes $i$ on a $\rm G$}-regular graph varies according to the transition rates
$\widetilde{T}_{i}^{+}(m,{\rm G},K)$ $=mK\frac{\widetilde{E}_{\rm G}(i)}{q_{\rm G}}\left[\rho_{M}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{\widetilde{E}_{\rm G}(i)}\right],$ (42)
$\widetilde{T}_{i}^{-}(m,{\rm G},K)$ $=mK\frac{\widetilde{E}_{\rm G}(i)}{q_{\rm G}}\left[\rho_{W}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{\widetilde{E}_{\rm G}(i)}\right].$
It is important to note that the $\widetilde{E}_{\rm G}(i)$ in these expressions explicitly represent the number of active edges between $M$ and $W$ on the metapopulation given $i$ mutant demes, taking into account the presence of extinct demes. Therefore, their expressions for a given graph structure $\rm G$ generally differ from those considered in the competition-dominated regime, denoted $E_{\rm G}(i)$. With these rates, we can solve the following first-step analysis equations for the probability $\phi_{{\rm int},i}^{\rm G}$ that the dynamical equilibrium comprising initially $i$ demes of type $M$ consists of occupied $M$ demes and extinct demes after a mean time $\theta_{{\rm int},i}^{\rm G}$:
@@
These equations are subject to the boundary conditions $\phi_{{\rm int},0}^{\rm G}=0,\phi_{{\rm int},\Omega_{\rm occ}}^{\rm G}=1$ and $\theta_{{\rm int},0}^{\rm G}=\theta_{{\rm int},\Omega_{\rm occ}}^{\rm G}=0$. We thus have $\phi_{\rm int}^{\rm G}\equiv p_{\text{surv}}\phi_{{\rm int},1}^{\rm G}$ and $\theta_{\rm int}^{\rm G}\equiv p_{\text{surv}}\theta_{{\rm int},1}^{\rm G}+(1-p_{\text{surv}})\tau_{E}$. The factor $p_{\text{surv}}=\frac{\Omega_{\text{occ}}}{\Omega}=1-1/\psi$ is the probability that the initial $M$ deme reaches the dynamical equilibrium (after what is assumed to be a short transient), while the contribution $(1-p_{\text{surv}})\tau_{E}$ to $\theta_{\rm int}^{\rm G}$ accounts for the probability that the initial $M$ deme goes extinct in a mean time $\tau_{E}$ (given by Eq. (7)) before reaching the equilibrium. The final expressions of $\phi_{\rm int}^{\rm G}$ and $\theta_{\rm int}^{\rm G}$ thus read
@@
$\widetilde{\gamma}_{\rm G}(i)\equiv\frac{\rho_{W}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{\widetilde{E}_{\rm G}(i)}}{\rho_{M}+\frac{1}{\psi-1}\frac{q_{\rm G}}{\Omega}\frac{i(\Omega_{\text{occ}}-i)}{\widetilde{E}_{\rm G}(i)}},$ (45)
@@
We notice that for all graphs $\rm G$}, the expressions of Eq. (44) coincide with those of Eq. (14) of the competition-dominated regime, with $\widetilde{\gamma}_{G}(i)\stackrel{{\scriptstyle\psi\gg 1}}{{\longrightarrow}}\gamma=\rho_{W}/\rho_{M}$. In Fig. 11, we find that the predictions of Eq. (44) are in good agreement with simulation results for all spatial structures $\rm G$}. Moreover, we observe that the spatial dependence of $\phi^{\rm G}_{\text{int}}$ and $\theta^{\rm G}_{\text{int}}$ is barely noticeable.
@@ Appendix G Fixation probability in time-switching environments under weak bottlenecks
In static environments, where $K$ is constant, a generalisation of the circulation theorem guarantees that the fixation probability is independent of the migration rate and spatial structure of the metapopulation arranged on a circulation graph; see Eq. (14). This results from a correspondence between the fixation probability and the number of $M$ demes performing a biased random walk on $\{0,\dots,\Omega\}$ with a bias that is independent of $m$ and $\rm G$} [41].
In time-switching environments under weak bottlenecks (deme extinction is negligible) the correspondence is between the fixation $\Phi^{\rm G}(\nu,m)$ and the random walk (with absorbing boundaries) on $\{0,\dots,\Omega\}\times\{-1,1\}$ for the number of fully mutant demes in the environmental state $\xi=\pm 1$. As a consequence, $\Phi^{\rm G}(\nu,m)$ is the probability of absorption in the state $\Omega$. In this setting, defining the state of the random walk by $(i,\xi)$, where $i=0,1,\dots,\Omega$, the random walk moves to the right ($i\to i+1$) with a probability $r(i,\xi)$, to the left ($i\to i-1$) with a probability $\ell(i,\xi)$, or switches environment ($\xi\to-\xi$) with probability $\epsilon(i,\xi)$, where
@@
$\Phi^{\rm G}(\nu,m)$ thus coincides with the probability that the random walk defined by Eq. (47) gets absorbed in the state $i=\Omega$. Unlike the case of the competition-dominated regime under the static environment, the fixation probability here typically depends on all parameters, including the migration rate. For the fixation probability to remain unchanged under parameter changes requires strict conditions. This can be seen by assuming that for a parameter set $S_{1}$ the probabilities $r_{1}(i,\xi)$, $\ell_{1}(i,\xi)$, and $\epsilon_{1}(i,\xi)$ define the random walk corresponding to the fixation probability $\Phi_{i,\xi}^{\rm G}$. We therefore have
$\Phi_{i,\xi}^{\rm G}=r_{1}(i,\xi)\Phi_{i+1,\xi}^{\rm G}+\ell_{1}(i,\xi)\Phi_{i-1,\xi}^{\rm G}+\epsilon_{1}(i,\xi)\Phi_{i,-\xi}^{\rm G}.$ (48)
We can also assume that under another set of parameters, say $S_{2}$, $\Phi_{i,\xi}^{\rm G}$ remains unchanged for all $i$ and $\xi$ with a corresponding random walk defined by the probabilities $r_{2}(i,\xi)$, $\ell_{2}(i,\xi)$, and $\epsilon_{2}(i,\xi)$, such that
$\Phi_{i,\xi}^{\rm G}=r_{2}(i,\xi)\Phi_{i+1,\xi}^{\rm G}+\ell_{2}(i,\xi)\Phi_{i-1,\xi}^{\rm G}+\epsilon_{2}(i,\xi)\Phi_{i,-\xi}^{\rm G}.$ (49)
Subtracting the second equation from the first, and using conservation of probability, we find that
$(r_{1}(i,\xi)-r_{2}(i,\xi))(\Phi_{i+1,\xi}^{\rm G}-\Phi_{i,-\xi}^{\rm G})$ (50)
$+(\ell_{1}(i,\xi)-\ell_{2}(i,\xi))(\Phi_{i-1,\xi}^{\rm G}-\Phi_{i,-\xi}^{\rm G})=0.$
The cases of $\Phi_{i+1,\xi}^{\rm G}-\Phi_{i,-\xi}^{\rm G}=0$ and $\Phi_{i-1,\xi}^{\rm G}-\Phi_{i,-\xi}^{\rm G}=0$ imply that the fixation probability of all transient states is identical, and therefore we neglect this unphysical case. Defining $\Delta r(i,\xi)=r_{1}(i,\xi)-r_{2}(i,\xi)$ and $\Delta\ell(i,\xi)=\ell_{1}(i,\xi)-\ell_{2}(i,\xi)$, Eq. (50) yields
${\Delta r(i,\xi)}=\frac{\Phi_{i+1,\xi}^{\rm G}-\Phi_{i,-\xi}^{\rm G}}{\Phi_{i,-\xi}^{\rm G}-\Phi_{i-1,\xi}^{\rm G}}{\Delta\ell(i,\xi)}.$ (51)
For each of the $2(\Omega-2)$ transient states we therefore have a constraint given by Eq. (51). However, $\Delta r(i,\xi)$ and $\Delta\ell(i,\xi)$ are controlled by $|S_{1}|=|S_{2}|=p$ degrees of freedom (system parameters) where $p\ll 2(\Omega-2)$. Therefore, the system is overdetermined and Eq. (51) is only generally satisfied across all $(i,\xi)$ for the trivial solution $\Delta r(i,\xi)=\Delta\ell(i,\xi)=0$, i.e. $S_{1}=S_{2}$. Thus, in time-fluctuating environments the fixation probability $\Phi^{\rm G}(\nu,m)$ is expected to depend on $m$ and $\rm G$}.
Interestingly however, Fig. 4(c) illustrates the almost unnoticeable dependence of $\Phi^{\rm G}(\nu,m)$ on the specific spatial structure. This is due to the overall similar impact of the factor $E_{\rm G}(i)/q_{G}$ for the various graphs. While differences arise when $m$ varies at fixed $\nu$ due to large variations in the timescales of the competition dynamics, varying spatial structure produces small changes in these timescales, and as such leads to only unnoticeable changes in $\Phi$.